# U-Nets as Belief Propagation: Efficient Classification, Denoising, and Diffusion in Generative Hierarchical Models

### v1

@@ 1 Introduction
The generative hierarchical model [Mos16, SFW24, TW24, PCT+23] discussed herein is a tree-structured graphical model, which has been widely employed in language and image generative modeling [Cho59, Lee96, AZL23, LGO00, Wil02, JG06]. We detail the precise definition of such generative hierarchical models in Section 2. The belief propagation denoising algorithm, which computes the denoising function in these models exactly, includes a downward process and an upward process, with the latter reusing the intermediate results from the downward process. In Section 4, we demonstrate how this algorithm naturally induces the encoder-decoder structure, the long skip connections, and the pooling and up-sampling operations of the U-Nets. This gives rise to an efficient sample complexity bound for learning the denoising function in generative hierarchical models using U-Nets. Additionally, we discuss the broader implications of these findings for diffusion models in generative hierarchical models.
@@ 3 The warm-up problem: Classification in GHMs
exhibiting a polynomial dependence on ${d}$ and $1/{\varepsilon}$. Such polynomial scaling aligns with existing literature [PMR+17, MSS20, SH20b, AZL22], which indicates that learning hierarchical models using multi-layer networks avoids the curse of dimensionality. Theorem 1 serves as a warm-up result in the classification context. In Section 4, we aim to extend similar methodologies to address denoising problems, employing analogous proof strategies.
@@ 3.1 Proof strategy: ConvNets approximate the belief propagation algorithm
In the following, we describe our strategy to control the approximation error: we first derive the belief propagation and message passing algorithm for computing the Bayes classifier ${\mu_{\star}}$, and then demonstrate that ConvNets are capable of effectively approximating this message passing algorithm.
@@ 4 Denoising and diffusion in GHMs
    $\hat{\bm{m}}_{t}\equiv\arg\min_{{\rm NN}_{t}\in{\mathcal{F}}}\frac{1}{n}\sum_{i=1}^{n}\|{\bm{x}}_{i}-{\rm NN}_{t}({\bm{x}}_{i}+t^{-1}{\bm{g}}_{i})\|_{2}^{2},~{}~{}~{}~{}{\bm{g}}_{i}\sim_{iid}{\mathcal{N}}({\bm{0}},{\mathbf{I}}_{d}).$ (ERM)
@@ 4.1 Proof strategy: U-Nets approximate the belief propagation algorithm
In the following, we describe our strategy to control the approximation error: we first derive the belief propagation and message passing algorithm for computing the Bayes denoiser ${{\bm{m}}_{\star}}$, and then demonstrate that U-Nets are capable of effectively approximating this message passing algorithm.
@@ 5 Further related work
Convolutional networks [LBD+89, LBBH98, KSH12, SLJ+15, HZRS16] have become the state-of-the-art architecture for image classification and have been the backbone for many computer vision tasks. U-Nets [RFB15, ZRSTL18, SPED21, OSF+18] have been particularly well-suited for image segmentation and denoising tasks [RFB15], and have served as the backbone architecture for diffusion models [SDWMG15, HJA20, SE19, SSDK+20]. A series of theoretical works has attempted to explain the inductive bias of CNNs [BM13, GLSS18, SH20a, LZA20, Bie21, MMM21, MM22, CFW23, FCW21, BVB21, Xia22, TW24, WW24]. However, they mostly focused on the classification and regression setting and were not concerned with the role of U-Nets in denoising tasks. The implicit bias of U-Nets has been theoretically investigated in [WFD+24, FWD+22], where they found that the U-Nets are conjugate to the ResNets. In contrast, we demonstrate that U-Nets can effectively approximate the belief propagation denoising algorithms of GHMs.
Generative hierarchical models.
Hierarchical modeling of data distributions has been proposed in a series of works [Mos16, PMR+17, MSS20, SH20b, AZL22, SFW24, TW24, PCT+23]. Recent studies [SFW24, TW24, PCT+23] empirically evaluated the performance of neural networks for classification and denoising tasks in generative hierarchical models. In particular, [PCT+23, TW24] have demonstrated that learning using deep convolutional networks can be significantly more efficient than employing two-layer fully-connected networks. The empirical evidence presented in [SFW24] indicates that generative hierarchical models are powerful tools for modeling combinatorial properties in image datasets.
@@
In recent years, diffusion models [SDWMG15, HJA20, SE19, SSDK+20] have emerged as a leading approach for generative modeling. Neural network-based score function approximation has been recently studied from the function approximation viewpoint in [OAS23, CHZW23, YHN+23, SCK23, BM23], and from the algorithm approximation viewpoint in [MW23a]. For a comprehensive introduction to the theory of diffusion models, see the recent review [CMFW24].

### v2

@@ 1 Introduction
The generative hierarchical model discussed herein is a tree-structured graphical model, which has been widely employed in language and image generative modeling [Cho59, Lee96, AZL23, LGO00, Wil02, JG06]. We detail the precise definition of such generative hierarchical models in Section 2. A series of recent work [Mos16, SFW24, TW24, PCT+23] have pioneered the use of generative hierarchical models in studying classification tasks and diffusion models. As noted by [SFW24], the belief propagation denoising algorithm, which computes the denoising function in these models exactly, includes a downward process and an upward process, with the latter reusing the intermediate results from the downward process. In Section 4, we demonstrate how this algorithm naturally induces the encoder-decoder structure, the long skip connections, and the pooling and up-sampling operations of the U-Nets. This gives rise to an efficient sample complexity bound for learning the denoising function in generative hierarchical models using U-Nets. Additionally, we discuss the broader implications of these findings for diffusion models in generative hierarchical models.
@@ 2 The generative hierarchical model
We remark that a series of studies [SFW24, TW24, PCT+23] have provided both theoretical and empirical evidence supporting the efficacy of GHMs as powerful tools for modeling the properties of combinatorial data.
@@ 3 The warm-up problem: Classification in GHMs
exhibiting a polynomial dependence on ${d}$ and $1/{\varepsilon}$. Such polynomial scaling aligns with existing literature [PMR+17, MSS20, SH20b, AZL22, PCT+23], which indicates that learning hierarchical models using multi-layer networks avoids the curse of dimensionality. Theorem 1 serves as a warm-up result in the classification context. In Section 4, we aim to extend similar methodologies to address denoising problems, employing analogous proof strategies.
@@ 3.1 Proof strategy: ConvNets approximate the belief propagation algorithm
In the following, we describe our strategy to control the approximation error: we first present the belief propagation and message passing algorithm for computing the Bayes classifier ${\mu_{\star}}$, and then demonstrate that ConvNets are capable of effectively approximating this message passing algorithm.
@@ 4 Denoising and diffusion in GHMs
    $\hat{\bm{m}}_{t}\equiv\arg\min_{{\rm NN}_{t}\in{\mathcal{F}}}\frac{1}{n{d}}\sum_{i=1}^{n}\|{\bm{x}}_{i}-{\rm NN}_{t}(t\cdot{\bm{x}}_{i}+\sqrt{t}\cdot{\bm{g}}_{i})\|_{2}^{2},~{}~{}~{}~{}{\bm{g}}_{i}\sim_{iid}{\mathcal{N}}({\bm{0}},{\mathbf{I}}_{d}).$ (ERM)
@@ 4.1 Proof strategy: U-Nets approximate the belief propagation algorithm
In the following, we describe our strategy to control the approximation error: we first present the belief propagation and message passing algorithm for computing the Bayes denoiser ${{\bm{m}}_{\star}}$, and then demonstrate that U-Nets are capable of effectively approximating this message passing algorithm.
@@
We remark that the downward-upward structure of belief propagation in generative hierarchical models has been pointed out in the literature [SFW24].
@@ 5 Further related work
Generative hierarchical models.
Hierarchical modeling of data distributions has been proposed in a series of works [Mos16, PMR+17, MSS20, SH20b, AZL22, PCT+23, SFW24, TW24]. While the hierarchical models in [PMR+17, MSS20, SH20b, AZL22] remain deterministic, [Mos16, PCT+23, SFW24, TW24] studied the generative version of hierarchical models. The theoretical and empirical evidence presented in [SFW24, TW24, PCT+23] underscores the effectiveness of generative hierarchical models in capturing the combinatorial properties of image datasets. Given their significant relevance to this work, we delve deeper into these studies.
Contributions of [PCT+23, SFW24, TW24].
The series of works on hierarchical generative models [PCT+23, SFW24, TW24] inspired the current study. [SFW24] first pointed out that the belief propagation denoising algorithm of hierarchical models consists of downward and upward processes. Through mean-field analysis on a random generative hierarchical model, they identified a phase transition phenomenon, aligning with empirical observations in diffusion models, thereby providing strong evidence of the efficacy of these models in handling combinatorial data properties. [PCT+23], on the other hand, first introduced these models in a classification context. [PCT+23, TW24] demonstrated that learning hierarchical models using multi-layer networks circumvents the curse of dimensionality. Specifically, they theoretically and empirically characterized the sample complexity, showing that it remains polynomial in dimension when learning convolutional networks under random generative rules. On the other hand, in the absence of correlations, they showed that the sample complexity is again exponential in the dimension, even for hierarchical generative models. This learning incapability is not captured by our analysis which does not consider optimization.
@@
Convolutional networks [LBD+89, LBBH98, KSH12, SLJ+15, HZRS16] have become the state-of-the-art architecture for image classification and have been the backbone for many computer vision tasks. U-Nets [RFB15, ZRSTL18, SPED21, OSF+18] have been particularly well-suited for image segmentation and denoising tasks [RFB15], and have served as the backbone architecture for diffusion models [SDWMG15, HJA20, SE19, SSDK+20]. A series of theoretical works has explained the inductive bias of CNNs [BM13, GLSS18, BM19b, BM19a, SH20a, LZA20, Bie21, MMM21, MM22, CFW23, FCW21, BVB21, Xia22, PCT+23, TW24, WW24]. However, they mostly focused on the classification and regression setting and were not concerned with the role of U-Nets in denoising tasks. The implicit bias of U-Nets has been theoretically investigated in [WFD+24, FWD+22], where they found that the U-Nets are conjugate to the ResNets. In contrast, we demonstrate that U-Nets can effectively approximate the belief propagation denoising algorithms of GHMs. We also note that [CKVEZ23] has analyzed the learning dynamics for a simple U-Net in diffusion models.
@@
In recent years, diffusion models [SDWMG15, HJA20, SE19, SSDK+20] have emerged as a leading approach for generative modeling. Neural network-based score function approximation has been recently studied from the function approximation viewpoint in [OAS23, CHZW23, YHN+23, SCK23, BM23], and from the algorithm approximation viewpoint in [MW23a]. Theoretical studies of other aspects of diffusion models include [LWYL22, LWCC23, LLT23, CCSW22, CDD23, CCL+22, CCL+23, CLL23, BDBDD23, EAMS22, MW23b, Cel22, GDKZ23, BM23, BBdBM24, CKVEZ23, FYWC24, WCL+24]. For a comprehensive introduction to the theory of diffusion models, see the recent review [CMFW24].